A 65 mm diameter solid shaft is to be welded to a flat plate and is required to carry a torque of 1750 N-m. If fillet joint is used for welding what will be the minimum size of the weld when working shear stress is 66 MPa.


1
Expert's answer
2021-03-18T03:36:25-0400

"= (T*d\/2)J"

where T = torsion

d = diameter of the shaft

J = polar moment of inertia

"= (T*d/2)/ (\Pi*t*d^3/4)""=2T/ (\Pi*t*d^2)"

t = s sin 45= 0.707s

"shear stress = 2T/ (\Pi * 0.0707s *d^2)"

"66* 10^6 = 2.83*1750/(\Pi * s *0.065^2)"

"1\/s = 176.88 M"

"s = 5.653 *10^-3M"

Size of the weld = 5.653 mm.